If X is a binomial random variable, for what value of Theta the probability is maximum?
1
Expert's answer
2019-10-17T09:18:37-0400
"X\sim bin(k;n,\theta)"

"f(k)=P(X=k)=\binom{n}{k}\theta^k(1-\theta)^{n-k}"

We can look at the ratio of successive outcomes


"r={P(X=k+1) \over P(X=k)}"

When "r>1" then "P(X=k+1)>P(X=k)"

When "r<1" then "P(X=k+1)<P(X=k)"

Maximum of the distribution occurs when "r" switches from being greater than "1" to less than "1."


"r={P(X=k+1) \over P(X=k)}={\binom{n}{k+1}\theta^{k+1}(1-\theta)^{n-(k+1)} \over \binom{n}{k}\theta^k(1-\theta)^{n-k}}="

"={n-k \over k+1}\cdot{\theta \over1-\theta}"

 What value of k results in "r\leq1?"


"{n-k \over k+1}\cdot{\theta \over1-\theta}\leq1"

"\theta(n-k)\leq(1-\theta)(k+1)"

"n\theta-k\theta\leq k+1-k\theta-\theta"

"k\geq n\theta-(1-\theta)"

 Max probability is the smallest interger value of "k\geq n\theta-(1-\theta)"


Likelihood function


"L(\theta|n, x)=\binom{n}{x}\theta^x(1-\theta)^{n-x}=""={n! \over x!(n-x)!}\theta^x(1-\theta)^{n-x}"

Log-likelihood function 


"\ln{L(\theta|n, x)}=\ln{\bigg({n! \over x!(n-x)!}\bigg)}+x\ln{\theta}+(n-x)\ln({1-\theta})"

"{d({L(\theta|n, x)}) \over dp}=0+x\cdot{1 \over \theta}+(n-x)\cdot(-{1 \over 1-\theta})=0"

"x(1-\theta)-(n-x)\theta=0"

"x-x\theta-n\theta+x\theta=0"

"x-n\theta=0"

"\theta={x \over n}"

The maximum likelihood estimate for "\theta" is just the average.